If there are 35 questions what is the most i can get wrong to get a 90%? if i miss 5 what is my percentage?
Answer by oberobic(2304)

x/35 = .9
x = .9*35
x = 31.5
That means you have to get at least 32 correct to achieve at least a 90%.
.
30/35 * 100 = 6/7 * 100 = 85.71%